# Discovering a Novel Dynamics Mechanism for Charmonium Scattering

### v1

@@ Abstract
As shown in Fig. 3, our model successfully reproduces the overall features of the di-$J/\psi$ invariant mass spectrum. This indicates that our proposed scattering mechanism between two charmonium states is reasonable, providing a novel approach to understanding hadron scattering, especially for fully heavy hadrons. Furthermore, both the lineshape of the black solid lines and the $\chi^{2}/d.o.f$ values in Table 1 suggest a slight preference for the $0^{++}$ configuration in the di-$J/\psi$ invariant mass spectrum. However, the $\chi^{2}/d.o.f$ values are not close to 1, primarily because we did not include channels with thresholds near 6.6 GeV and 7.3 GeV, leading to mismatches in these regions. Despite this, our primary goal of validating the new scattering mechanism is achieved through the correspondence with experimental results.
In addition, Fig. 3 reveals a significant twist at 6.865 GeV, spanning about 10 MeV, in the partial contributions of some channels, such as the $J/\psi\psi(3686)$ channel in the $0^{++}$ fit. This is attributed to an anomalous threshold effect caused by $\psi(3770)$. Since $\psi(3770)$ is above the $D\bar{D}$ threshold, it can decay into a pair of real $D$ mesons when it interacts with another vector charmonium, as illustrated in Fig. 3. When $\psi_{3}=J/\psi$ and the remaining intermediate meson in the triangle loop is also a $D$ meson, the anomalous threshold effect occurs, resulting in the observed twist around 6.9 GeV. This kind of threshold effect has been extensively discussed in previous studies Lu:2023aal ; Zhuang:2021pci , and our results are consistent with Ref. Lu:2023aal . The presence of such an effect is a crucial feature of our model. More precise future experiments around 6.9 GeV would greatly aid in validating our model.
@@
Moreover, the wide peak around 6.6 GeV can be reproduced by interference effects without introducing additional channels. The sharp resonance-like structure in $\sigma_{J/\psi J/\psi}$ in both configurations suggests a potential $J/\psi J/\psi$ bound state, as it is not reflected in the overall spectrum due to phase space limitations. This conclusion aligns with Refs. Dong:2020nwy ; Dong:2021lkh ; Liang:2021fzr . Importantly, the dip near 6.8 GeV, particularly evident in the $J/\psi\psi(3686)$ contribution in the $2^{++}$ fit, indicates the possible existence of a state similar to $\Upsilon(10753)$, which also shows as a dip in $e^{+}e^{-}\to b\bar{b}$ cross section Dong:2020tdw . This pole around 6.8 GeV is also identified in Refs. Dong:2020nwy ; Liang:2021fzr ; Guo:2020pvt ; Zhuang:2021pci . Thus, detailed future experimental analyses of this dip could further test our model.
@@
Overall, our model not only reproduces the experimental data but also integrates nearly all conclusions from previous studies into a unified framework. This initial success in understanding the scattering mechanism between two charmonium states motivates further study and broader application of our model. This letter represents the first step in that direction.
@@
Summary.—Recently, the CMS Collaboration reported a high-precision measurement of the di-$J/\psi$ invariant mass spectrum CMS:2023owd , which confirmed the previous observation of $X(6900)$ by the LHCb Collaboration LHCb:2020bwg and revealed additional enhancements. The ATLAS Collaboration has also contributed to this effort, reinforcing these findings ATLAS:2023bft . These observations have spurred a variety of theoretical interpretations, including the existence of exotic hadronic states and dynamics generated by coupled channels liu:2020eha ; Tiwari:2021tmz ; Lu:2020cns ; Faustov:2020qfm ; Zhang:2020xtb ; Li:2021ygk ; Liu:2021rtn ; Giron:2020wpx ; Karliner:2020dta ; Wang:2020ols ; Ke:2021iyh ; Zhu:2020xni ; Jin:2020jfc ; Yang:2021hrb ; Albuquerque:2020hio ; Albuquerque:2021erv ; Wu:2022qwd ; Asadi:2021ids ; Yang:2020wkh ; Chen:2020xwe ; Deng:2020iqw ; Gordillo:2020sgc ; Zhao:2020zjh ; Faustov:2021hjs ; Wang:2021kfv . However, a critical issue remains unresolved: the nature of charmonium scattering. Traditional one-boson exchange mechanisms, highly effective for describing hadron scattering, fall short in quantitatively characterizing charmonium scattering. This shortcoming is particularly pronounced in fully heavy hadron systems.
@@
In this letter, we introduce a novel mechanism for charmonium scattering. Our approach leverages the internal structure of charmonium, where charm and anti-charm quarks can capture light flavor anti-quarks and quarks from the vacuum, forming virtual charmed mesons. These virtual mesons act as mediators, enabling an effective one-boson exchange process for charmonium scattering.
@@
Our proposed mechanism successfully reproduces the di-$J/\psi$ invariant mass spectrum data from CMS and LHCb CMS:2023owd ; LHCb:2020bwg , validating its effectiveness. A significant advantage of this approach is its broad applicability. It is not limited to charmonium scattering but can be extended to investigate the scattering of all fully heavy hadrons, an area of increasing interest within the physics community, including lattice QCD research groups Lyu:2021qsh ; Mathur:2022ovu . This approach addresses the limitations of traditional one-boson exchange mechanisms, which are inadequate for fully heavy hadron systems due to the absence of light quark exchanges. By incorporating virtual meson formation, our model provides a robust framework for understanding charmonium interactions.
This study provides a comprehensive framework for charmonium scattering, integrating well with existing experimental observations CMS:2023owd ; LHCb:2020bwg and theoretical efforts on scattering mechanisms Wang:2020wrp ; Wang:2020tpt ; Wang:2022jmb ; Gong:2020bmg ; Dong:2020nwy ; Dong:2021lkh ; Liang:2021fzr ; Lu:2023aal ; Guo:2020pvt ; Zhuang:2021pci . The success of our mechanism in explaining the di-$J/\psi$ spectrum sets a precedent for future studies on fully heavy hadron scattering. We emphasize the importance of further experimental investigations, particularly high-precision measurements around key energy thresholds, to refine and validate our model. Our work lays the groundwork for a deeper understanding of scattering mechanisms in fully heavy hadron systems, with potential implications for broader applications in particle physics.
@@
Acknowledgements.—This work is supported by the National Natural Science Foundation of China under Grant Nos. 12335001, 12247101, 12305139, and 12305087, the National Key Research and Development Program of China under Contract No. 2020YFA0406400, and the 111 Project under Grant No. B20063, the Fundamental Research Funds for the Central Universities, and the project for top-notch innovative talents of Gansu province. R. C. is supported by the Xiaoxiang Scholars Programme of Hunan Normal University. Q. H. is supported by the Start-up Funds of Nanjing Normal University under Grant No. 184080H201B20.

### v2

@@ II Application to reproduce the CMS and LHCb data
As shown in Fig. 3, our model successfully reproduces the overall features of the di-$J/\psi$ invariant mass spectrum. This indicates that our proposed scattering mechanism between two charmonium states is reasonable, providing a new approach to understanding hadron scattering, especially for fully heavy hadrons. Furthermore, both the lineshape of the black solid lines and the $\chi^{2}/d.o.f$ values in Table 1 suggest a slight preference for the $0^{++}$ configuration in the di-$J/\psi$ invariant mass spectrum, but $2^{++}$ configuration is also possible. Here, the $\chi^{2}/d.o.f$ values are not close to 1, primarily because there are no thresholds near 6.6 GeV and 7.3 GeV in our fit, leading to mismatches in these regions. Despite this, our primary goal of validating the new scattering mechanism is achieved through the correspondence with experimental results.
In addition, Fig. 3 reveals a twist at 6.865 GeV, which is just located at the $J/\psi\psi(3770)$ threshold. Obviously, from the diagram given in Fig. 2, it can be inferred that such phenomenon is attributed to the $\psi(3770)$, whose mass is above the $D\bar{D}$ threshold and thus can decay into a pair of real $D$ mesons. As a result, when the scattering energy reach up to the $J/\psi\psi(3770)$ threshold, rescattering between coupled channels exceed the branch point located at the $J/\psi\psi(3770)$ threshold, which can cause a cusp at this threshold Wang:2020wrp . In addition, since $\psi(3770)$ can be a real particle at this time, when the two charmed mesons linked with it are $D\bar{D}$, the branch point located at $D\bar{D}$ threshold is simultaneously triggered, which may result into further enhancements and generate a tall but very narrow peak just around the $J/\psi\psi(3770)$ threshold, and consequently only slightly modifies the shape of the peak near 6.9 GeV. As for the triangle singularity, based on the algorithm we used in our previous work Huang:2020kxf , it can occur when the intermediate states in the triangle loop is $D^{\ast}\bar{D}D$, with $D^{\ast}\bar{D}$ being linked with charmonia, and $D^{\ast}D$ connected to the light exchanged boson. However, at this time, the kinematical limitation needs the charmonia should be an intermediate virtual particle with energy in a range from 3.872 to 3.875 GeV, and the exchanged light boson should have energy less than 0.14 GeV. Apparently, such an effect is highly possible to be suppressed, since either the charmonia or the exchanged light boson is deviated from the center value of mass. Thus, the presence of such a strong threshold effect is a crucial feature of our model. More precise future experimental measurements around 6.9 GeV in the future would greatly aid in validating our model.
@@
Moreover, the wide peak around 6.6 GeV can be reproduced by interference effects without introducing additional channels. The sharp resonance-like structure in $\sigma_{J/\psi J/\psi}$ in both configurations suggests a potential $J/\psi J/\psi$ bound state, as it is not reflected in the overall spectrum due to phase space limitations. This conclusion aligns with Refs. Dong:2020nwy ; Dong:2021lkh ; Liang:2021fzr . Importantly, if seeing Table 2 and Fig. 3 together, the dip near 6.8 GeV, particularly evident in the $J/\psi\psi(3686)$ contribution in the $2^{++}$ fit, indicates the existence of a state similar to $\Upsilon(10753)$, which also shows as a dip in $e^{+}e^{-}\to b\bar{b}$ cross section Dong:2020tdw . This pole around 6.8 GeV is also identified in Refs. Dong:2020nwy ; Liang:2021fzr ; Guo:2020pvt ; Zhuang:2021pci . Thus, detailed future experimental analyses of this dip could further test our model.
@@
Overall, it indicates that for either $0^{++}$ or $2^{++}$ configurations, there always exists a states below the $J/\psi J/\psi$ threshold. In addition, $X(6600)$ is not a real physical state but an effect of interference, while the dip around 6.8 GeV correspond to a pole mainly generated by the $J/\psi\psi(3686)$ channel. The peak around 6.9 GeV is dominantly generated by a physical state emerging from the $J/\psi\psi(3770)$ channel, but the strong threshold cusp effect there will also make an explicit twist on the shape of the peak. Due to the reason that our mechanism integrates nearly all the conclusions from previous studies into a unified framework, this initial success in understanding the scattering mechanism between two charmonium states motivates further study and broader application of our model. This work represents the first step in that direction.
III Summary
@@
Recently, the CMS Collaboration reported a high-precision measurement of the di-$J/\psi$ invariant mass spectrum CMS:2023owd , which confirmed the previous observation of $X(6900)$ by the LHCb Collaboration LHCb:2020bwg and revealed additional enhancements. The ATLAS Collaboration has also contributed to this effort, reinforcing these findings ATLAS:2023bft . These observations have spurred a variety of theoretical interpretations, including the existence of exotic hadronic states and dynamics generated by coupled channels liu:2020eha ; Tiwari:2021tmz ; Lu:2020cns ; Faustov:2020qfm ; Zhang:2020xtb ; Li:2021ygk ; Liu:2021rtn ; Giron:2020wpx ; Karliner:2020dta ; Wang:2020ols ; Ke:2021iyh ; Zhu:2020xni ; Jin:2020jfc ; Yang:2021hrb ; Albuquerque:2020hio ; Albuquerque:2021erv ; Wu:2022qwd ; Asadi:2021ids ; Yang:2020wkh ; Chen:2020xwe ; Deng:2020iqw ; Gordillo:2020sgc ; Zhao:2020zjh ; Faustov:2021hjs ; Wang:2021kfv . However, a critical issue remains unresolved: the nature of charmonium scattering. Traditional one-boson exchange mechanisms, highly effective for describing hadron scattering, fall short in quantitatively characterizing charmonium scattering. This shortcoming is particularly pronounced in fully heavy hadron systems.
@@
To solve this puzzle, in this work, we introduce a new mechanism for charmonium scattering. Our approach leverages the internal structure of charmonium, where charm and anti-charm quarks can capture light flavor anti-quarks and quarks from the vacuum, forming virtual charmed mesons. These virtual mesons act as mediators, enabling an effective one-boson exchange process for charmonium scattering.
@@
Our proposed mechanism successfully reproduces the di-$J/\psi$ invariant mass spectrum data from CMS and LHCb CMS:2023owd ; LHCb:2020bwg , validating its effectiveness. A significant advantage of this approach is its broad applicability. It is not limited to charmonium scattering but can be extended to investigate the scattering of all fully heavy hadrons, an area of increasing interest within the physics community, including lattice QCD research groups Lyu:2021qsh ; Mathur:2022ovu . With our new mechanism, we can easily predict that there should also exist several states in the di-$\chi_{cJ}$ system, which are all very closed to their corresponding thresholds. Furthermore, due to the heavy quark symmetry, our mechanism should also exist in $\Upsilon\Upsilon$, $J/\psi\Upsilon$, $\chi_{bJ}\chi_{bJ^{\prime}}$, $\chi_{cJ}\chi_{bJ^{\prime}}$ channels, which can result into the generations of near threshold states. Nevertheless, this approach addresses the limitations of traditional one-boson exchange mechanisms, which are inadequate for fully heavy hadron systems due to the absence of light quark exchanges. It lays the groundwork for a deeper understanding of scattering mechanisms in fully heavy hadron systems, with potential implications for broader applications in particle physics. Thus, We emphasize the importance of further experimental investigations, particularly high-precision measurements around key energy thresholds, to refine and validate our model.
IV Note added
We noticed that the CMS collaboration has measured the spin-parity of all-charm tetraquarks CMS:2025ecr . The result showed that the quantum numbers of these states prefer $J^{PC}=2^{++}$, which corroborates the conclusions of this work.
Acknowledgements
@@
This work is supported by the National Natural Science Foundation of China under Grant Nos. 12335001, 12247101, 12305139, 12305087, and 12475080, the National Key Research and Development Program of China under Contract No. 2020YFA0406400, the ‘111 Center’ under Grant No. B20063, the Natural Science Foundation of Gansu Province (No. 22JR5RA389, No. 25JRRA799), the fundamental Research Funds for the Central Universities (No. lzujbky-2023-stlt01), the project for top-notch innovative talents of Gansu province, Lanzhou City High-Level Talent Funding, and the Xiaoxiang Scholars Programme of Hunan Normal University.